Rio Bus Tours has incurred the following bus maintenance costs during the recent tourist season. (The real is Brazil’s national monetary unit. On the day this exercise was written, the real was equivalent in value to 0.2545 U.S. dollar.)

MonthMiles Traveled by Tour BusesCost
November
15,450

18,900real
December
17,700

19,200
January
19,950

19,350
February
24,300

19,800
March
31,800

21,000
April
13,800

18,300


Required:

 1. Use the high-low method to estimate the variable cost per tour mile traveled and the fixed cost per month.

 2. Develop a formula to express the cost behavior exhibited by the company's maintenance cost.

 3. Predict the level of maintenance cost that would be incurred during a month when 47,000 tour miles are driven.
Answer #1

1.

Highest activity cost = 21,000 real

Lowest activity cost = 18,300 real

Highest activity = 31,800

Lowest activity = 13,800

Variable cost per mile = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (21,000-18,300)/(31,800-13,800)

= 2,700/18,000

= 0.15 real per mile

Fixed cost = Highest activity cost - Highest activity x Variable cost per miles

= 21,000- 31,800 x 0.15

= 21,000-4,770

= 16,230 real

2.

Cost Fourmula

y = 16,230+ 0.15 X

Total cost = Fixed cost + Variable cost per mile x Activity miles driven

3.

Total Cost ,when 47,000 tour miles are driven= Fixed cost + Variable cost per mile x Activity miles driven

= 16,230+ 0.15 x 47,000

= 16,230+7,050

= 23,280 real